### v2

# CastMind: An Interaction-Driven Agentic Reasoning Framework for Cognition-Inspired Time Series Forecasting

Abstract
Time series forecasting plays a crucial role in decision-making across many real-world applications. Despite substantial progress, most existing methods still treat forecasting as a static, single-pass regression problem. In contrast, human experts form predictions through iterative reasoning that integrates temporal features, domain knowledge, case-based references, and supplementary context, with continuous refinement. In this work, we propose CastMind, an interaction-driven agentic reasoning framework that enables accurate time series forecasting with training-free large language models. CastMind reformulates forecasting as an expert-like process and organizes it into a multi-stage workflow involving context preparation, reasoning-based generation, and reflective evaluation, transforming forecasting from a single-pass output into a multi-turn, autonomous interaction process. To support diverse perspectives commonly considered by human experts, we develop a lightweight toolkit comprising a feature set, a knowledge base, a case library, and a contextual pool that provides external support for LLM-based reasoning. Extensive experiments across multiple benchmarks show that CastMind generally outperforms representative baselines [^1] https://github.com/SkyeGT/CastMind.
Machine Learning, ICML
1 Introduction
Figure 1: Illustration of the motivation behind CastMind: shifting forecasting from static, single-pass regression to human-like, interactive, multi-turn reasoning.
Time series forecasting plays a crucial role in decision-making across various real-world applications, such as energy (Cao et al., 2023), climate (Wang et al., ), and industry (Ma et al., 2025). Despite advances with statistical (Webby & O’Connor, 1996), deep learning (Benidis et al., 2022), and recent foundation models (Das et al., 2024), these methods typically treat forecasting as a static, single-pass regression task, generating predictions based solely on historical observations (Brockwell & Davis, 2002). This approach neglects reasoning, interaction, and iterative refinement, which are critical in real-world decision-making.
In contrast, human experts rarely follow this single-pass model. Instead, forecasting is seen as an iterative reasoning process, where predictions are continuously and gradually refined by integrating diverse information. Experts consider external context, domain knowledge, and similar past cases (Lawrence et al., 2006), revising predictions through repeated analysis and self-consistency checks. This expert-driven process is multi-turn and interactive, prioritizing reasoning and refinement over direct input-output mapping.
While existing forecasting models are effective, they struggle to capture this expert-like reasoning. Statistical and deep learning methods (Chen et al., ; Ma et al., 2025) focus on predictive accuracy through fixed input-output mappings, which are limited in modeling reasoning and self-refinement. Recent advances in foundation models (Das et al., 2024; Liu et al., 2025b) and large language models (LLMs) (Chang et al., 2023; Liu et al., 2024b; Zhang et al., 2025) extend modeling capacity, yet they are often applied to time series forecasting in a single-pass generation manner or as black-box predictors (Zhou et al., 2023; Jin et al., 2023) , lacking mechanisms for context exploration, hypothesis adjustment, or reflective verification. This leaves a gap between existing models and the iterative, interaction-driven processes observed in expert practice, as illustrated in Figure 1.
To address this, we argue that time series forecasting be reformulated as an interaction-driven agentic reasoning problem. Rather than a single-pass prediction, forecasting becomes a multi-turn process where an agent actively prepares contextual information, reasons over diverse sources, and refines predictions through self-reflection. To operationalize this perspective, We introduce CastMind, an interaction-driven agentic reasoning framework for time series forecasting. CastMind leverages training-free LLMs as reasoning engines and organizes forecasting into a structured workflow: context preparation, reasoning-based generation, and reflective verification. By modeling interaction and refinement, CastMind bridges the gap between existing models and expert-like forecasting processes.
CastMind implements this through a congnitive-inspired workflow that separates reasoning from prediction. It starts with context preparation, gathering relevant information beyond raw time series, including temporal features, domain knowledge, case-based references and supplementary context. The LLM then analyzes this context and produces a intermediate forecast through reasoning. A reflective verification stage follows, ensuring consistency and plausibility, enabling self-correction when necessary. To support this process, CastMind uses a lightweight toolkit with core components: a feature set, knowledge base, case library,and contextual pool. Together, the workflow and toolkit enable CastMind to conduct forecasting as an interactive, multi-turn reasoning process that closely mirrors expert practice.
The contributions of this work are as follows:
  • We reformulate time series forecasting from a static, single-pass regression task into an interaction-driven agentic reasoning problem, highlighting the role of iterative reasoning and refinement in forecasting.
  • We propose CastMind, a training-free agentic reasoning framework that integrates a structured workflow with a lightweight forecasting toolkit for context-aware and reflective forecasting.
  • Extensive experiments across multiple benchmarks show CastMind consistently outperforms statistical, deep learning, and foundation-model baselines, demonstrating its effectiveness.
2 Related Work
2.1 Time Series Forecasting
Time series forecasting has evolved from manual expert analysis, often based on visual inspection, to automated statistical methods and deep learning approaches (Qiu et al., 2024). Early heuristic approaches lacked scalability, driving the adoption of statistical models like ARIMA (Hyndman & Khandakar, 2008) and ETS (Gardner Jr, 1985). While formal and interpretable, these methods struggle with nonlinearities and high-dimensional dynamics. Subsequently, deep learning architectures (Li et al., 2025; Huang et al., ; Wang et al., 2024b), particularly Transformers (Chen et al., ; Ilbert et al., 2024; Liu et al., 2023), became standard in real-world applications by effectively capturing long-term dependencies and complex patterns (Challu et al., 2023; Woo et al., 2022; Götz et al., 2024). Recently, foundation models such as Sundial (Liu et al., 2025b) and Chronos (Ansari et al., 2024) have demonstrated strong generalization capabilities across diverse tasks via large-scale pretraining. Despite their zero-shot performance, most remain static, single-pass predictors. They lack the contextual reasoning, reflection, and interactive capabilities with external tools that are critical for human-level forecasting in dynamic environments.
2.2 LLM-based Reasoning
Recent advancements in LLMs have demonstrated remarkable generalization capabilities in reasoning and decision-making (Guo et al., 2025). Foundational is in-context learning (ICL), enabling models to adapt to unseen tasks via prompts without parameter updates (Brown et al., 2020; Xue & Salim, 2023). To elicit multi-step inference, prompting strategies like chain-of-thought (CoT) (Wei et al., 2022) decompose complex problems into intermediate reasoning steps. Subsequent research enhanced robustness through aggregation like self-consistency (Wang et al., 2022) and decomposition like least-to-most prompting (Zhou et al., 2022). Moving beyond static reasoning, a significant paradigm shift models LLMs as autonomous agents operating in multi-step, goal-directed settings. Unlike passive inference, agentic systems maintain intermediate states and interleave reasoning with actions, invoking tools to ground outputs. Frameworks like ReAct (Yao et al., 2023), Toolformer (Schick et al., 2023), and WebGPT (Nakano et al., 2021) exemplify this approach by integrating search engines and APIs to improve reliability. To handle increasingly complex tasks, recent agents evolved into comprehensive cognitive architectures. These systems extend tool use by incorporating modules like structured planning (Wang et al., 2023; Erdogan et al., 2025), long-term memory (Packer et al., 2023), and self-correction like reflexion (Shinn et al., 2023). This modular design enables refinement and adaptive execution, forming the basis for our framework.
Figure 2: Overall architecture of CastMind. The three-stage process (Investigator, Generator, Reflector) bridges human wisdom and LLM intelligence, simulating expert cognition to transform forecasting into an interactive, multi-turn reasoning process.
3 The Proposed CastMind
In this section, we introduce CastMind, which reformulates forecasting as a multi-stage process that simulates human-like reasoning to prepare, generate, and evaluate predictions.
3.1 Problem Definition
Let $X_{\text{en}}\in\mathbb{R}^{H}$ denote the endogenous time series over a look-back window of length $H$. Let $X_{\text{ex}}\in\mathbb{R}^{(d-1)\times(H+L)}$ denote the (d1)(d-1) exogenous variables, where $X_{\text{ex}}^{(j)}\in\mathbb{R}^{H+L}$ represents the $j$-th exogenous variable of length $H+L$. In practical scenarios, experts can leverage future exogenous data, such as weather forecasts, to make more accurate predictions. Incorporating this known future data allows the framework to better capture external information, improving forecasting accuracy.
3.2 Overall Agentic Workflow
As shown in Figure 2, CastMind is a multi-stage interaction-driven agentic reasoning framework designed for cognition-inspired time series forecasting. It reformulates forecasting from a static, single-pass mapping task into an interactive, multi-turn iterative process with training-free LLM reasoning. Before the workflow begins, CastMind prepares a specialized toolkit comprising a feature set, a knowledge base, a case library, a contextual pool. The process then unfolds in three coordinated stages. Initially, CastMind organizes useful contextual input and information retrieved from the toolkit to construct a well-defined forecasting context. Subsequently, the framework operates on this foundation to generate reasoning-based predictions. Finally, it evaluates intermediate outputs and the reason trace to detect inconsistencies and iteratively refine the predictions. Overall, this workflow mirrors how human experts extract context, reason about predictions, and adjust decisions.
3.3 Forecasting Toolkit Design
We design a lightweight toolkit to support reasoning with multi-source information, mirroring how experts look beyond raw historical observations.
Feature Set.
CastMind extracts temporal features from endogenous time series and exogenous variables to summarize temporal dynamics. Given a group of extraction functions $\mathcal{F}=\{f_{1},\dots,f_{n}\}$, the feature set $F$ is constructed by applying each function to all input series:
$F=\{f(X)\mid f\in\mathcal{F},\;X\in\{X_{\text{en}},\dots,X_{\text{ex}}^{(d-1)}\}\},$ (1)
which aggregates a set of twenty temporal features. Definitions of all features are provided in Appendix A.1.
Knowledge Base.
CastMind maintains a knowledge base that provides prior information to support reasoning and decision-making. The knowledge base contains: (1) practical rules, which are experience-based guidelines derived from empirical observations. These rules capture common patterns in the data, such as the effect of rainfall on water levels or the correlation between hot weather and increased electricity demand. (2) theoretical principles, which are domain-specific laws and theorems that help constrain the model’s predictions. These include well-established principles such as Joule’s law and the conservation of mass, which ensure that the model’s predictions align with physical and scientific laws. Let KKdenote the domain knowledge retrieved from the knowledge base.
Case Library.
To simulate the expert use of historical experiences, CastMind constructs a case library by associating each training sample $\langle X_{i},Y_{i}\rangle$ with its optimal candidate model $M_{i}$ selected from the candidate pool $\mathcal{M}$ (detailed in Appendix B.2). To enable efficient retrieval and generation of case-based predictions, we perform K-means clustering (detailed in Appendix A.2) on all $X_{i}$, grouping them into multiple clusters. Each cluster $C_{m}$ contains training samples $X_{i}$ that are most similar to one another. The center $c_{m}$ of each cluster $C_{m}$ is determined by averaging all the samples within the cluster, where the center is the mean of the samples $X_{i}$ in $C_{m}$, calculated as $c_{m}=\frac{1}{|C_{m}|}\sum_{X_{i}\in C_{m}}X_{i}$.
Contextual Pool.
In forecasting, human experts often incorporate contextual information to support predictions. The contextual pool provides supplementary context, such as holidays, to enhance the framework’s understanding. Denoted as $S$, this context is supplied in textual form, complementing historical data for more accurate forecasting under complex conditions where numerical patterns alone are insufficient (Liu et al., 2025a). Importantly, all supplementary context is external to the dataset, preventing any data leakage that could influence the forecasting task.
3.4 The Investigator: Context Preparation
Accurate forecasting requires a well-defined prediction task, a proper understanding of data semantics, and the availability of relevant contextual information. Therefore, CastMind constructs a standardized contextual input $I_{\text{in}}$ that integrates time series with textual descriptions. Specifically, $I_{\text{in}}$ comprises the task prompt, the data profile (detailed meanings of the variables), original timestamp information (Wang et al., 2024a; Zeng et al., 2024), and the raw time series $(X_{\text{ex}},X_{\text{en}})$ in original space. This standardized input serves as the foundation for subsequent analysis.
Upon receiving $I_{\text{in}}$, the investigator analyzes the task to identify the most relevant temporal features, denoted as $F_{\text{sel}}$, from the feature set. The nearest cluster center $c_{m}$ to the input $X_{\text{en}}$ is then determined using Euclidean distance. For the case-based prediction $Y_{\text{case}}$, the outputs from the models in the selected cluster $C_{m}$ are averaged, as shown by
$Y_{\text{case}}=\frac{1}{|C_{m}|}\sum_{j\in C_{m}}M_{j}(X_{\text{en}}).$ (2)
At the same time, the framework retrieves the nearest neighbor $X_{\text{nb}}$ and its corresponding prediction $Y_{\text{nb}}$, offering both aggregate trends and specific historical references to inform the reasoning process. The framework then aggregates the input, selected features, and auxiliary resources from the toolkit, including domain knowledge $K$, supplementary context $S$, and case-based references $Y_{\text{case}}$, $X_{\text{nb}}$, and $Y_{\text{nb}}$, into a comprehensive forecasting context:
$I_{\text{con}}=(I_{\text{in}},F_{\text{sel}},K,S,Y_{\text{case}},X_{\text{nb}},Y_{\text{nb}}),$ (3)
where consolidated context $I_{\text{con}}$ is then passed to the next stage, providing a reliable foundation for generative reasoning and reflective refinement.
3.5 The Generator: Reasoning-based Forecasting
Given the consolidated context $I_{\text{con}}$, CastMind derives an intermediate prediction $Y_{\text{int}}$ through reasoning-driven generation instead of single-pass regression. The model initializes with the case-based prediction $Y_{\text{case}}$ as a reliable reference. To handle distribution shifts between neighbor $X_{\text{nb}}$ and the current sequence $X_{\text{en}}$, CastMind uses the neighbor trajectory $Y_{\text{nb}}$ to adjust for scale and fluctuation differences (Liu et al., 2025c). The framework also incorporates temporal features from the selected feature set $F_{\text{sel}}$, along with domain knowledge $K$ and supplementary context $S$. These elements support more sophisticated adjustments, such as shifting or reshaping, that allow the model to adapt to subtle but important changes in the data, which may be overlooked in simpler models. By incorporating timestamps $T$ for time-aware calibration (e.g., distinguishing weekday variations), the reasoning process is formalized as $Y_{\text{int}}=\mathcal{G}(I_{\text{con}})$, where $\mathcal{G}$ denotes the generator that integrates this context to output the adjusted $Y_{\text{int}}$ for the subsequent stage.
Table 1: MSE and MAE results on short-term and long-term forecasting benchmarks (lower is better; best results are highlighted in bold, and second best results are underlined). WP and SP denote the Windy Power and Sunny Power datasets, respectively.
Setting Short-term Forecasting Long-term Forecasting
Metric Model BE DE FR NP PJM ETTh ETTm WP SP MOPEX
MSE Sundial 651.237 264.618 942.081 28.667 30.704 9.437 3.211 2167.943 100.304 5.149
Chronos 625.634 223.153 803.076 22.180 25.695 9.397 2.749 2268.519 79.245 5.283
DLinear 658.530 239.928 811.453 32.215 42.154 8.506 2.631 1932.200 19.058 5.256
PatchTST 627.149 208.888 797.263 24.634 31.874 8.396 2.622 2269.641 19.435 5.358
TimesNet 636.660 209.366 929.100 32.116 34.890 7.940 2.439 1930.876 18.426 4.955
TimeXer 702.862 252.001 834.693 27.306 25.876 8.765 2.557 2169.818 20.114 4.825
iTransformer 606.528 229.955 940.227 27.088 35.131 8.307 3.136 2063.985 17.956 4.883
Autoformer 890.843 331.306 934.364 46.097 77.570 11.598 4.224 2567.613 39.129 6.168
Prophet 992.900 320.631 1035.704 57.794 52.261 46.697 21.183 8071.244 64.807 12.395
SNaive 857.119 415.723 915.798 44.893 41.387 10.753 2.746 1948.056 79.611 7.180
ARIMA 869.444 372.910 963.997 55.507 33.021 10.879 2.709 2166.644 80.478 9.190
CES 808.608 315.576 1084.713 40.776 38.327 24.218 4.124 9364.980 78.713 6.399
CastMind 536.454 193.829 721.023 27.113 25.151 7.641 2.414 1548.825 13.294 4.771
MAE Sundial 10.997 10.707 8.536 3.362 3.987 2.314 1.258 31.218 6.687 1.422
Chronos 9.623 9.999 7.536 3.176 3.630 2.250 1.180 29.617 4.566 1.307
DLinear 12.833 12.833 10.449 3.842 4.645 2.239 1.138 29.325 2.735 1.333
PatchTST 11.435 10.042 9.126 3.266 4.166 2.169 1.114 32.462 2.782 1.439
TimesNet 11.300 10.197 11.099 3.805 4.342 2.079 1.066 29.398 2.781 1.350
TimeXer 10.700 10.259 9.377 3.436 3.651 2.204 1.113 31.914 2.915 1.277
iTransformer 11.242 10.227 11.093 3.434 4.373 2.136 1.277 29.893 2.673 1.337
Autoformer 16.619 13.329 13.362 4.856 6.426 2.637 1.571 37.002 4.339 1.958
Prophet 16.942 13.254 14.422 4.999 5.533 4.592 3.113 56.059 6.442 2.503
SNaive 13.704 13.362 11.155 4.102 4.648 2.469 1.186 27.876 4.558 1.619
ARIMA 17.038 11.816 13.224 4.445 4.004 2.429 1.168 28.767 4.598 1.852
CES 14.414 11.451 12.144 3.993 4.467 3.646 1.483 42.040 4.555 1.470
CastMind 9.612 9.925 7.133 3.145 3.596 2.017 1.057 24.540 1.843 1.353
3.6 The Reflector: Iterative Refinement
In real-world forecasting, predictions are usually reviewed and refined before being finalized. Accordingly, CastMind includes a reflection stage to verify the prediction and support iterative refinement. This stage improves the overall reliability and reproducibility of the framework and consists of two parts: (1) prediction evaluation, which assesses whether the prediction satisfies the task requirements and is supported by the available evidence, ensuring that the prediction is reasonable and reliable; and (2) reasoning-trace evaluation, which examines the generated reasoning trace to ensure that each step is valid and consistent, and identifies potential issues such as unsupported claims or logical gaps. Together, these checks help ensure that both the forecast and its underlying reasoning adhere to the task requirements. If issues are found, CastMind applies targeted, evidence-supported corrections to $Y_{\text{int}}$ to obtain a refined prediction. The refined result is then re-evaluated with the same criteria. If acceptable, CastMind outputs the forecast; otherwise, it returns feedback to the previous stage for another update, repeating until an acceptable result is obtained.
4 Experiments
In this section, we conduct experiments on various forecasting benchmarks comparing CastMind with mainstream approaches to demonstrate its effectiveness.
4.1 Experimental Setup
Datasets.
We evaluate our framework on a diverse set of benchmarks covering both short-term and long-term horizons with rich contextual dependencies. For short-term forecasting, we employ five regional datasets from the EPF benchmark (Wang et al., 2024c), which track electricity prices alongside exogenous variables such as generation and load forecasts. In the long-term setting, the ETTh and ETTm datasets (Zhou et al., 2021) are utilized to monitor transformer temperature and load variations. To capture renewable energy dynamics, we adopt the Windy and Sunny Power datasets (iFLYTEK AI Challenge, 2025), which align power generation records with meteorological conditions. Additionally, the MOPEX dataset (Schaake et al., 2006) is included for hydrological forecasting, featuring streamflow series supported by climatic factors. Detailed dataset descriptions are provided in Appendix B.1.
Baselines.
We include twelve time series forecasting models covering three representative categories: (1) statistical models: Prophet (Taylor & Letham, 2018), SNaive (Hyndman & Athanasopoulos, 2018), ARIMA (Hyndman & Khandakar, 2008), and CES (Svetunkov et al., 2022); (2) deep learning models: DLinear (Zeng et al., 2023), PatchTST (Nie et al., 2022), TimesNet (Wu et al., 2022), TimeXer (Wang et al., 2024c), iTransformer (Liu et al., 2023), and Autoformer (Wu et al., 2021); and (3) foundation models: Sundial (Liu et al., 2025b) and Chronos (Ansari et al., 2024). These representative baselines comprehensively validate our framework’s effectiveness.
Implementation Details.
Statistical models are implemented using standard formulations, whereas deep learning models are developed with the TSLib (Wang et al., 2024b) toolkit under the official settings. Foundation models are evaluated using their publicly released implementations. For short-term forecasting, we adopt the standard protocol of N-BEATSx (Olivares et al., 2023), where the input length is 168 and the horizon is 24. For long-term forecasting, we fix the look-back window at 96 and the prediction length at 96.
Table 2: Comprehensive ablation study evaluating the impact of the Retrieval Toolkit, Reasoning, and Reflective Validation. “w/o Toolkit” denotes the removal of all retrieved components, “w/o Reasoning” corresponds to the best performance of representative candidate models (serving as the backbone baseline), and “w/o Reflexion” indicates the exclusion of the self-correction mechanism.
Model BE FR NP Sunny Power ETTh ETTm
MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
w/o Toolkit 685.412 13.058 812.556 9.942 42.109 4.675 55.321 4.218 14.882 3.105 5.211 1.844
w/o Reasoning 606.528 11.242 797.263 9.126 22.180 3.176 17.956 2.673 7.940 2.079 2.439 1.066
w/o Reflexion 551.302 11.536 741.347 8.791 28.739 3.359 40.184 3.387 10.952 2.841 4.013 1.576
Full Model 536.454 9.612 721.023 7.133 27.113 3.145 13.294 1.843 7.641 2.017 2.414 1.057
Figure 3: Component-wise ablation analysis of the forecasting context. The chart illustrates the sensitivity of model performance to the removal of specific information.
4.2 Main Results
Table 1 compares CastMind against twelve baselines across short-term and long-term benchmarks. Overall, CastMind achieves state-of-the-art performance with superior accuracy across most metrics. In short-term scenarios, CastMind shows outstanding stability in the face of high volatility, effectively managing unpredictable fluctuations. Unlike statistical methods such as ARIMA, which struggle with rapid non-linear fluctuations and inherent stochastic volatility in price dynamics, or deep learning models such as PatchTST, which may over-smooth dynamics, CastMind effectively captures sharp peaks by leveraging historical references to avoid phase shifts. This gives it a significant advantage on datasets like BE and FR. In long-term settings, CastMind exhibits superior forecasting capability, especially on datasets like Windy Power and Sunny Power, which exhibit complex environmental dependencies. While foundation models such as Chronos perform competitively due to their zero-shot generalization, CastMind maintains a consistent lead through its reasoning and refinement capabilities that enforce causal constraints derived from diverse contexts. For instance, on the Windy Power dataset, CastMind substantially outperforms baselines like DLinear. This superiority across different scales validates the effectiveness of the interactive, multi-turn reasoning process in handling diverse temporal patterns and ensuring forecasting stability.
Table 3: Ablation study evaluating the contribution of each component in the proposed forecasting toolkit to overall forecasting performance, using MSE as the evaluation metric.
Datasets BE DE Windy Power ETTm
w/o Feature Library 624.267 221.970 2442.103 3.595
w/o Knowledge Base 641.524 211.768 2579.324 4.486
w/o Case Library 607.572 253.103 1670.633 3.900
w/o Contextual Pool 613.258 231.134 2001.323 3.653
Full Model 536.454 193.829 1548.825 2.414
4.3 Ablation studies
This section validates the CastMind framework through a multi-perspective ablation study, covering both architectural stages, contextual input, and auxiliary toolkit modules.
Ablation Study on Framework Components.
To validate the workflow, we conduct an ablation study by removing the toolkit, reasoning, and reflexion. As shown in Table 2, the full model consistently achieves best performance, validating component collaboration. Pecifically, removing the toolkit yields highest errors in scenarios like BE and ETTm, confirming external context is foundational. Furthermore, comparing the baseline lacking reasoning, which serves as the state-of-the-art candidate model backbone, with the full model reveals that reasoning-based forecasting substantially improves accuracy, evidenced by notable error reduction on the FR dataset. Most importantly, results on the Sunny Power dataset highlight the indispensability of the reflexion module, where the variant excluding self-correction performs worse than even the non-reasoning baseline. This anomaly suggests that without self-correction, generative reasoning may be prone to hallucinations or overfitting, an issue rectified by the full model acting as a safeguard for reliability and physical consistency.
Figure 4: Comparison of forecasting performance across different LLM backbones on multiple datasets.
Ablation Study on Contextual Input.
To verify the necessity of constructing a comprehensive context $I_{\text{con}}$, we conduct a component-wise ablation study. As shown in Figure 3, the complete CastMind framework consistently yields the lowest MAE, confirming the synergistic contribution of all elements. Specifically, removing timestamps causes significant degradation on periodic datasets such as PJM and Sunny Power, highlighting the critical role of explicit temporal identifiers in capturing seasonality that numerical sequences alone cannot convey. Similarly, the exclusion of exogenous variables and static data profile leads to substantial error increases on tasks like ETTm and NP, demonstrating that external drivers and domain-specific inductive biases are essential for reasoning about complex system behaviors beyond historical autoregressive patterns.
Figure 5: Impact of the number of clusters. The optimal granularity balances matching specificity with consensus stability.
Ablation Study on Forecasting Toolkit
To verify the contribution of each auxiliary module, we conduct a component-wise ablation study. As presented in Table 3, the full model consistently outperforms all variants, validating the value of the complete toolkit. Notably, removing the knowledge base causes drastic deterioration on physics-governed datasets such as Windy Power and ETTm, indicating external domain knowledge is essential for reasoning about complex dynamics which numerical history cannot reveal. Similarly, excluding the feature set and case library leads to a marked decline on volatile datasets like BE and DE, confirming their role in providing references to stabilize predictions against noise. Furthermore, the contextual pool acts as a vital buffer, with its removal consistently degrading performance by disrupting the understanding of time series evolution.
4.4 Exploration Analysis
Figure 6: Feature-level and model-level case analysis of CastMind across datasets. The heatmaps report the relative usage frequency (%) of temporal features and candidate forecasting models, highlighting a preference for statistical features and deep learning models.
Figure 7: Case-study comparisons on four datasets (ETTm, MOPEX, Windy Power, and DE), where CastMind produces predictions that track the ground truth more closely than the cased-based prediction, especially under periodic or highly volatile patterns.
Table 4: Impact of generation strategies. Holistic generation outperforms the step-wise baseline, demonstrating the necessity of coherent, one-shot prediction.
Datasets BE PJM Windy Power
Metric MSE MAE MSE MAE MSE MAE
Step-wise Generation 616.026 10.881 31.872 4.102 2164.368 29.829
Holistic Generation (Ours) 536.454 9.612 25.151 3.596 1548.825 24.540
Impact of Forecasting Backbone.
To investigate the impact of the foundation LLM on predictive performance, we evaluate the framework using diverse LLM backbones, including DeepSeek-V3 (Liu et al., 2024a), DeepSeek-R1 (Guo et al., 2025), GPT-4o (Achiam et al., 2023), Gemini-2.5-pro (Comanici et al., 2025), and GPT-5 (Singh et al., 2025). As illustrated in Figure 4, GPT-5 consistently achieves the lowest MAE across diverse datasets such as Windy Power, DE, and PJM, establishing a clear performance advantage over other LLMs. The superior performance of GPT-5 is attributed to its powerful pre-training and reasoning capabilities, enabling it to effectively handle complex time series, adapt to various context, and provide more accurate predictions. Notably, the performance gap becomes most distinct on the Windy Power dataset, where GPT-5 demonstrates a significantly stronger ability to model complex environmental dependencies and high-frequency volatility compared to competitive baselines.
Impact of Cluster Number.
To investigate the impact of historical scenario organization, we evaluate forecasting performance across varying cluster counts. As illustrated in Figure 5, the error exhibits a concave pattern, initially improving then degrading as the number of clusters increases. Specifically, with a small number of clusters, the library suffers from over-generalization, indiscriminately grouping heterogeneous look-back windows together. Consequently, the synthesized prediction, derived from diverse expert models, becomes a diluted generic signal lacking specificity. Conversely, increasing the cluster count beyond the optimal range leads to excessive fragmentation, resulting in sparse clusters. This sparsity undermines consensus stability, making guidance sensitive to noise. Thus, moderate granularity is crucial for ensuring the accuracy of the prediction.
Impact of Generation Strategy.
To determine the optimal generation strategy for time series forecasting with LLMs, we conduct an ablation study comparing our default holistic generation against a step-wise generation baseline. In the baseline setting, the model is explicitly prompted to generate the forecast horizon in sequential segments rather than outputting the entire trajectory in a single coherent stream. As shown in Table 4, the step-wise approach results in a significant performance degradation across all datasets. For instance, on the Windy Power, the error increases drastically compared to the holistic approach. This deterioration suggests that forcing the model to fragment the prediction process disrupts the continuity of temporal dependencies. Holistic generation serves as a superior strategy by maintaining a global receptive field, allowing the LLM to plan the long-term trend and seasonality consistently without being interrupted by artificial segmentation boundaries.
4.5 Case Studies
Preference Analysis on Toolkit.
We analyze the usage frequency within the feature set and case library to understand CastMind’s preferences. As shown in Figure 6, heatmaps reveal selection patterns driven by data characteristics. In the feature domain, CastMind favors fundamental statistical features, suggesting that reliable global properties serve as better foundations than volatile complexity measures. Meanwhile, model selection shows a clear bias towards deep learning architectures, with Transformers consistently dominating the candidate pool over statistical models, confirming that high-capacity backbones provide the necessary prior. Ultimately, the varying distribution across datasets confirms that the framework dynamically adjusts its retrieval strategy, rather than relying on static priors.
Case-based Prediction Analysis.
To assess forecasting quality, we compare final predictions with the case-based prediction and ground truth. As shown in Figure 7, the case-based prediction often suffers from over-smoothing, generating conservative trajectories that miss rapid fluctuations. This is evident on MOPEX and Windy Power, where the candidate models misses significant peaks and reverts to the mean. In contrast, CastMind shows superior sensitivity, closely tracking the ground truth even under high volatility on the ETTm and DE. This confirms that reasoning allows the model to escape the mean-reversion trap, ensuring forecasts retain the original series’ characteristics. A detailed case study is provided in the AppendixB.4.
5 Conclusion
In this work, we introduces CastMind, an interaction-driven agentic reasoning framework designed to bridge the gap between existing time series forecasting models and expert-driven iterative reasoning. CastMind redefines the forecasting process as a multi-turn iteration, integrating context preparation, reasoning-based generation, and reflective evaluation, transforming forecasting from a single-pass output into a multi-turn autonomous interaction process. The proposed framework leverages training-free LLMs as reasoning engines, complemented by a lightweight toolkit that effectively combines human wisdom with LLM intelligence. Comprehensive evaluations demonstrate that CastMind’s ability to model expert-like iterative reasoning and refinement is key to its superior performance in both short-term and long-term forecasting tasks. Although the results are preliminary, our findings suggest that forecasting is shifting from a static regression-based approach to a interaction-driven agentic reasoning paradigm.
Impact Statement
This work advances time series forecasting through CastMind, an interaction-driven agentic framework that enhances transparency and decision reliability by mimicking expert-like iterative reasoning. By transforming forecasting into a multi-stage, autonomous interaction process supported by external context, our approach significantly improves interpretability in critical sectors such as infrastructure management and financial planning. However, we acknowledge that the multi-turn interaction with LLMs incurs higher inference latency and energy consumption compared to static, single-pass regression models. Consequently, we emphasize that deployment in high-stakes environments should remain subject to rigorous human oversight to mitigate potential risks associated with autonomous agentic decision-making. The information in the contextual pool and knowledge base is not directly sourced from the dataset itself, thus there is no risk of data leakage.
Appendix A Method Details
A.1 Feature Overview
As detailed in Table 5, the temporal feature set consists of 20 distinct metrics that collectively capture both distributional properties and sequential dependencies of the time series. These features include moment-based metrics, such as mean, variance, skewness, and kurtosis, which provide a global overview of the series’ shape and scale. Additionally, temporal features capture the sequential structure and dynamics of the time series. For example, autocorrelation-based metrics (e.g., acf1, acf10) quantify linear correlations (Box et al., 2015), while spectral entropy assesses the signal complexity and regularity (Hyndman et al., 2015). To capture structural heterogeneity, we integrate metrics such as lumpiness, flat spots, and crossing points, which help identify state transitions and geometric characteristics (Kang et al., 2017). Lastly, seasonal strength measures the dominance of seasonal patterns relative to residual variation (Wang et al., 2006), guiding the model to adaptively handle strong or weak seasonality.Collectively, this comprehensive temporal feature set equips CastMind with a deeper understanding of the time series, supporting effective reasoning and self-correction.
Table 5: Mathematical definitions of the twenty extracted temporal features used in CastMind, covering distributional statistics, temporal dependence, spectral complexity, structural heterogeneity, and seasonality, serving as a compact summary of key temporal characteristics for forecasting.
Feature Description Formula / Mathematical Definition
basic_count Sample size $N=|\{x_{t}\}|$
basic_mean Arithmetic Mean $\mu=\frac{1}{N}\sum_{t=1}^{N}x_{t}$
basic_std Standard Deviation $\sigma=\sqrt{\frac{1}{N-1}\sum_{t=1}^{N}(x_{t}-\mu)^{2}}$
basic_min/max Extremes $\min(x_{t}),\quad\max(x_{t})$
basic_skew Skewness (Fisher-Pearson) $\text{Skew}=\frac{1}{N}\sum_{t=1}^{N}\left(\frac{x_{t}-\mu}{\sigma}\right)^{3}$
basic_kurt Excess Kurtosis $\text{Kurt}=\left[\frac{1}{N}\sum_{t=1}^{N}\left(\frac{x_{t}-\mu}{\sigma}\right)^{4}\right]-3$
acf1 Autocorrelation at lag 1 $\rho_{1}=\frac{\sum_{t=1}^{N-1}(x_{t}-\mu)(x_{t+1}-\mu)}{\sum_{t=1}^{N}(x_{t}-\mu)^{2}}$
acf10 Sum of squares of first 10 ACFs $\text{ACF}_{10}=\sum_{k=1}^{10}\rho_{k}^{2}$
diff1_acf1 ACF1 of first difference Let $y_{t}=x_{t}-x_{t-1}$, compute $\rho_{1}(y_{t})$
diff2_acf1 ACF1 of second difference Let $z_{t}=y_{t}-y_{t-1}$, compute $\rho_{1}(z_{t})$
spectral_entropy Shannon entropy of Normalized PSD (Welch) $H_{sp}=-\sum_{f}P(f)\log_{2}P(f)$, where $P(f)$ is the normalized PSD est. by Welch method.
entropy Spectral Entropy (Periodogram) Similar to above, but $P(f)$ estimated via standard Periodogram. Normalized by $\log(\text{length})$.
lumpiness Variance of variances on tiles Let $\sigma^{2}_{i}$ be the variance of the $i$-th non-overlapping window. $\text{Lump}=\text{Var}(\{\sigma^{2}_{1},\sigma^{2}_{2},\dots,\sigma^{2}_{M}\})$
flat_spots Max run length of discretized series $\max_{k}(\text{length of }k\text{-th consecutive run in equi-width bins})$
crossing_points Median crossing count $\sum_{t=1}^{N-1}\mathbb{I}((x_{t}-\text{med})(x_{t+1}-\text{med})<0)$
seasonal_strength Strength of seasonality (STL) $F_{s}=\max\left(0,1-\frac{\text{Var}(R_{t})}{\text{Var}(S_{t}+R_{t})}\right)$, based on STL decomp $x_{t}=T_{t}+S_{t}+R_{t}$.
season_acf1 Seasonal ACF $\rho_{S}$ where $S$ is the seasonal period (e.g., 24).
A.2 K-means Clustering
K-means is a classical unsupervised clustering method that partitions a set of samples into $K$ clusters by minimizing the within-cluster sum of squared distances. Given a dataset $\mathcal{X}=\{x_{i}\}_{i=1}^{N}$ with $x_{i}\in\mathbb{R}^{d}$ and a predefined number of clusters $K$, K-means aims to learn a set of centroids $\{\mu_{k}\}_{k=1}^{K}$ and an assignment function $c(\cdot)$ that maps each sample to a cluster index.
Objective.
K-means minimizes the following objective:
$\min_{\{\mu_{k}\}_{k=1}^{K},\,c}\sum_{i=1}^{N}\left\lVert x_{i}-\mu_{c(i)}\right\rVert_{2}^{2},$ (4)
where $c(i)\in\{1,\dots,K\}$ denotes the cluster assignment of $x_{i}$ and $\mu_{k}$ denotes the centroid of cluster $k$.
Algorithm.
K-means alternates between an assignment step and an update step until convergence:
Algorithm 1 K-means Clustering
0: Dataset $\mathcal{X}=\{x_{i}\}_{i=1}^{N}$, number of clusters $K$
0: Cluster assignments $\{c(i)\}_{i=1}^{N}$ and centroids $\{\mu_{k}\}_{k=1}^{K}$
1: Initialize centroids $\{\mu_{k}\}_{k=1}^{K}$ (e.g., random initialization or K-means++)
2:repeat
3:  for each sample $x_{i}$ do
4:   Assign cluster index
$c(i)\leftarrow\arg\min_{k\in\{1,\dots,K\}}\lVert x_{i}-\mu_{k}\rVert_{2}^{2}$
5:  end for
6:  for each cluster $k=1,\dots,K$ do
7:   Update centroid
$\mu_{k}\leftarrow\frac{1}{|\mathcal{C}_{k}|}\sum_{i:c(i)=k}x_{i}$
8:  end for
9:until cluster assignments converge or objective change is below a threshold
Practical notes.
K-means is sensitive to initialization and may converge to a local optimum. In practice, we run K-means multiple times with different initializations and choose the solution with the lowest objective value. The per-iteration complexity is $\mathcal{O}(NKd)$, where $N$ is the number of samples, $K$ is the number of clusters, and $d$ is the feature dimension.
Appendix B Experimental Details
B.1 Datasets
We conduct short-term forecasting on electricity price forecasting datasets (Lago et al., 2021), which include five datasets representing different day-ahead electricity markets, each with a sampling frequency of 1 hour. The descriptions of the datasets are as follows: (1) BE represents Belgium’s electricity market, which records hourly electricity prices, Belgium’s load forecasts, and generation forecasts from France. (2) DE represents the German electricity market, which records hourly electricity prices, zonal load forecasts in the Amprion TSO area, and wind and solar power generation forecasts. (3) FR represents the electricity market in France, containing hourly electricity prices, along with corresponding load and generation forecasts. (4) NP represents the Nord Pool electricity market, which records hourly electricity prices and provides corresponding grid load and wind power forecasts. (5) PJM represents the Pennsylvania-New Jersey-Maryland market, which includes zonal electricity prices for the Commonwealth Edison (COMED) area, along with system load and COMED load forecasts.
For long-term forecasting, we used five datasets, and the names and basic descriptions of these datasets are as follows: (1) ETTh1 contains oil temperature data and load features from six electricity transformers, with a sampling frequency of 1 hour. (1) ETTm1 contains oil temperature data and load features from six electricity transformers, with a sampling frequency of 15 minutes. (3)Windy Power dataset comes from the iFLYTEK AI Developer Competition and represents the actual power generation from a wind farm, along with six types of meteorological data: Direct Radiation, Wind Direction at 80m, Wind Speed at 80m, Temperature at 2m, Relative Humidity at 2m, and Precipitation. Its sampling frequency is 15 minutes. (4)Sundy Power dataset also comes from the iFLYTEK AI Developer Competition and represents the actual power generation from another wind farm, including the same six meteorological data points as the Windy Power dataset, with a sampling frequency of 15 minutes.(5) MOPEX dataset contains time series data from the field of hydrology, providing daily streamflow discharge and four types of meteorological data: Mean Areal Precipitation (MAP), Climatic Potential Evaporation (CPE), Daily Maximum Air Temperature (Tmax), and Daily Minimum Air Temperature (Tmin). Its sampling frequency is 1 day. See Table 6 for the dataset description.
Table 6: Dataset descriptions. Ex. and En. are abbreviations for the Exogenous variable and Endogenous variable, respectively. #Num represents the number of exogenous variables in each dataset. The dataset size is organized in (Train, Validation, Test).
Dataset #Num Ex. Descriptions En. Descriptions Sampling Frequency Dataset Size
BE 2 Generation, System Load Belgium’s Electricity Price 1 Hour (10224,1584,3024)
DE 2 Wind power, Ampirion zonal load German’s Electricity Price 1 Hour (10224,1584,3024)
FR 2 Generation, System Load France’s Electricity Price 1 Hour (10224,1584,3024)
NP 2 Grid Load, Wind Power Nord Pool Electricity Price 1 Hour (10224,1584,3024)
PJM 2 System Load, SyZonal COMED load Pennsylvania-New Jersey-Maryland Electricity Price 1 Hour (10224,1584,3024)
ETTh1 6 Power Load Feature Oil Temperature 1 Hour (8544,1344,2544)
ETTm1 6 Power Load Feature Oil Temperature 15 Minutes (16896,2496,4896)
Windy Power 6 Meteorological Variables Power Generation 15 Minutes (16896,2496,4896)
Sunny power 6 Meteorological Variables Power Generation 15 Minutes (16896,2496,4896)
MOPEX 4 Meteorological Variables Streamflow Discharge 1 day (8544,1344,2544)
B.2 Preliminary Model
We briefly summarize each forecasting baseline in the same order as introduced in the main text.
  • Prophet (Taylor & Letham, 2018). Prophet is a decomposable forecasting model that combines trend, seasonality, and event/holiday effects within an additive formulation. It is designed to be robust to missing data and outliers and supports flexible seasonality modeling.
  • Seasonal Naive (SNaive) (Hyndman & Athanasopoulos, 2018). SNaive is a simple seasonal benchmark that predicts future values by repeating the most recent observation from the same seasonal position (e.g., the same hour/day/month in the previous cycle). It provides a strong reference for highly seasonal series.
  • ARIMA (Hyndman & Khandakar, 2008). ARIMA models a time series using autoregressive and moving-average terms, combined with differencing to handle non-stationarity. It remains a standard statistical baseline for univariate forecasting with interpretable components.
  • CES (Svetunkov et al., 2022). Complex exponential smoothing extends classical exponential smoothing by leveraging complex-valued representations to better capture certain temporal structures. It is a lightweight statistical method that can serve as a competitive baseline for a range of series types.
  • DLinear (Zeng et al., 2023). DLinear is a simple linear forecasting model that applies trend–seasonality decomposition and uses linear projections to extrapolate each component. Despite its simplicity, it is a strong long-horizon forecasting baseline.
  • PatchTST (Nie et al., 2022). PatchTST segments time series into patches that are treated as tokens for a Transformer encoder, improving efficiency and enabling longer-context modeling. It further adopts channel-independent processing to enhance scalability for multivariate forecasting.
  • TimesNet (Wu et al., 2022). TimesNet models multi-periodic temporal variations by transforming 1D sequences into 2D representations based on discovered periods. It uses a parameter-efficient backbone to capture intra-period and inter-period patterns and has shown strong performance across forecasting tasks.
  • iTransformer (Liu et al., 2023). iTransformer inverts the tokenization perspective by treating variables as tokens and applying attention across variates to better capture multivariate correlations. This design improves effectiveness for long look-back windows while retaining the standard Transformer components.
  • Autoformer (Wu et al., 2021). Autoformer integrates series decomposition into the network and replaces standard self-attention with an autocorrelation mechanism to discover long-range periodic dependencies at the sub-series level. It is a representative decomposition-based Transformer for long-term forecasting.
  • Sundial (Liu et al., 2025b). Sundial is a generative time series foundation model pre-trained on large-scale time series corpora, aiming to support strong zero-shot generalization. It produces probabilistic forecasts by modeling the distribution of future patches rather than only point predictions.
  • Chronos (Ansari et al., 2024). Chronos reframes time series forecasting as a language-modeling problem by scaling and quantizing values into discrete tokens and training Transformer language models on these sequences. Forecasts are obtained by sampling future token trajectories, enabling probabilistic and often competitive zero-shot forecasting.
B.3 Complete Experiment
CastMind utilizes a Case Library comprising various statistical and deep learning models to support its reasoning process. To provide a granular view of how our framework interacts with and improves upon these individual predictors, we report the full breakdown of experimental results in this appendix. The table 7 contrast the standalone performance of each individual candidate model available in the library against the final forecasting output of the CastMind framework. The results clearly indicate a synergistic effect: CastMind successfully mitigates the limitations of individual small models by introducing contextual reasoning and iterative refinement, yielding superior forecasting outcomes compared to the preliminary models alone.
Table 7: Forecasting performance on short-term and long-term benchmarks. We compare CastMind with each individual candidate model in the case library to show the benefit of interaction-driven reasoning and iterative refinement.
Setting Short-term Forecasting Long-term Forecasting
Models Metrics BE DE FR NP PJM ETTh ETTm Windy Power Sunny Power MOPEX
CastMind MSE 536.454 193.829 721.023 27.113 25.151 7.641 2.414 1548.825 13.294 4.771
MAE 9.612 9.925 7.133 3.145 3.596 2.017 1.057 24.540 1.843 1.353
Sundial MSE 651.237 264.618 942.081 28.667 30.704 9.437 3.211 2167.943 100.304 5.149
MAE 10.997 10.707 8.536 3.362 3.987 2.314 1.258 31.218 6.687 1.422
Chronos MSE 625.634 223.153 803.076 22.180 25.695 9.397 2.749 2268.519 79.245 5.283
MAE 9.623 9.999 7.536 3.176 3.630 2.250 1.180 29.617 4.566 1.307
DLinear MSE 658.530 239.928 811.453 32.215 42.154 8.506 2.631 1932.200 19.058 5.256
MAE 12.833 12.833 10.449 3.842 4.645 2.239 1.138 29.325 2.735 1.333
PatchTST MSE 627.149 208.888 797.263 24.634 31.874 8.396 2.622 2269.641 19.435 5.358
MAE 11.435 10.042 9.126 3.266 4.166 2.169 1.114 32.462 2.782 1.439
TimesNet MSE 636.660 209.366 929.100 32.116 34.890 7.940 2.439 1930.876 18.426 4.955
MAE 11.300 10.197 11.099 3.805 4.342 2.079 1.066 29.398 2.781 1.350
iTransformer MSE 606.528 229.955 940.227 27.088 35.131 8.307 3.136 2063.985 17.956 4.883
MAE 11.242 10.227 11.093 3.434 4.373 2.136 1.277 29.893 2.673 1.337
Autoformer MSE 890.843 331.306 934.364 46.097 77.570 11.598 4.224 2567.613 39.129 6.168
MAE 16.619 13.329 13.362 4.856 6.426 2.637 1.571 37.002 4.339 1.958
Prophet MSE 992.900 320.631 1035.704 57.794 52.261 46.697 21.183 8071.244 64.807 12.395
MAE 16.942 13.254 14.422 4.999 5.533 4.592 3.113 56.059 6.442 2.503
SNaive MSE 857.119 415.723 915.798 44.893 41.387 10.753 2.746 1948.056 79.611 7.180
MAE 13.704 13.362 11.155 4.102 4.648 2.469 1.186 27.876 4.558 1.619
ARIMA MSE 869.444 372.910 963.997 55.507 33.021 10.879 2.709 2166.644 80.478 9.190
MAE 17.038 11.816 13.224 4.445 4.004 2.429 1.168 28.767 4.598 1.852
CES MSE 808.608 315.576 1084.713 40.776 38.327 24.218 4.124 9364.980 78.713 6.399
MAE 14.414 11.451 12.144 3.993 4.467 3.646 1.483 42.040 4.555 1.470
CrostonClassic MSE 927.984 294.443 1020.099 47.462 99.401 9.719 3.287 2206.911 84.788 5.322
MAE 17.597 12.983 15.158 4.804 7.819 2.331 1.365 32.036 8.250 1.449
Optimizers MSE 750.780 424.497 855.155 68.113 36.066 11.772 3.533 2277.535 78.985 8.064
MAE 12.839 12.532 11.204 4.782 4.304 2.541 1.261 29.878 4.558 1.777
HistoricAverage MSE 745.450 381.713 955.659 54.432 108.607 10.309 3.298 2671.839 59.759 6.142
MAE 16.741 14.568 14.634 5.225 8.101 2.571 1.306 36.529 6.130 1.709
B.4 Case Studies of CastMind Refinement Process
CastMind utilizes a Case Library comprising statistical models, deep learning models, and foundation models. To provide a granular view of how CastMind interacts with and improves upon the case-based predictions produced by these candidiate models, we report the full breakdown of experimental results in this appendix. Table 7 contrasts the standalone performance of each individual model in the library with the final forecasting output produced by the CastMind framework. The results clearly indicate a synergistic effect: by introducing contextual reasoning and iterative refinement, CastMind mitigates the limitations of individual models and generally achieves superior forecasting performance compared to the preliminary models alone.
Case Study Analysis on Mopex Background Information: User Request: Submit a forecasting task to predict the ’Streamflow Discharge’ for the next 96 steps. Context: The historical data is highly volatile, but case-based prediction is currently a very smooth curve with small fluctuations, missing potential flood peaks. Investigator: Response: The candidiate model smooths out the output, likely to minimize MSE loss. I must ensure that no critical drivers are being overlooked. I need to scan the future ’Precipitation’ forecast to detect any extreme weather inputs. Tool Call: Scan for future exogenous features, focusing on the ’precipitation’ variable to detect values beyond the ’extreme_event’ threshold. Tool Call Result: Three distinct high-intensity rainfall spikes were detected at $t=15$, $t=50$, and $t=85$. Generator: Response: The Investigator Agent detected heavy rainfall, but the candidiate model predicts steady flow. This suggests that the rain is ’disappearing,’ which violates the Principle of Mass Conservation. If the input (rainfall) exceeds infiltration, it must generate output (runoff). Tool Call: Query the knowledge base regarding the ’rainfall-runoff mass balance principle.’ Tool Call Result: The returned principle is the ’Mass Conservation’ rule, which implies that the input mass (precipitation) cannot vanish. Significant rainfall input must result in a discharge spike (runoff). Response: Mass conservation must be enforced. I will generate an intermediate prediction (X_int) by injecting impulse spikes into the flow forecast, strictly corresponding to the rainfall timestamps. Tool Call: Generate intermediate prediction. Tool Call Result: X_int was generated with spikes added at $t=15$, $t=50$, and $t=85$. Reflector: Response: Reviewing X_int, the discharge peaks are perfectly synchronized with the rainfall peaks (zero lag). In a real catchment, water requires time to travel to the outlet (Time of Concentration). This violates temporal logic. Evaluation Result: FAIL. Reason: Unrealistic zero lag. The discharge peak must lag behind the rainfall peak due to the hydrological routing process. Refinement: A refinement is required. I will apply a ’Temporal Shift’ to the peaks (delaying by $k$ steps) and adjust the falling limb to follow a natural recession decay curve to match the physical routing behavior. Final Output: The final forecast (X_final) was generated, and after being evaluated by the Reflector Agent, it was deemed satisfactory and passed.
Case Study Analysis on ETTm Background Information: User Request: Submit a forecasting task to predict the ‘Oil Temperature‘ for the next 96 steps. Context: The historical target sequence is extremely stable, and the case-based prediction is a flat horizontal prediction. Investigator: Response: The candidiate model predicts a flat trend due to historical stability. However, it is necessary to verify whether the future system state remains stable. To this end, the future values of the primary driver ‘Power Load‘ are scanned to detect potential changes in operating conditions. Tool Call: Scan future exogenous features, focusing on the ‘power_load‘ variable over a 96-step horizon to detect pattern changes. Tool Call Result: Future power load exhibits a sharp V-shaped drop at $t=45$ (shutdown event), followed by a significant ramp-up starting at $t=60$. Generator: Response: The Investigator detects significant fluctuations in power load, while the candidiate model predicts a constant temperature. This contradicts physical causality. Therefore, the physical relationship between electrical load and transformer oil temperature must be retrieved. Tool Call: Query the knowledge base for the causal relationship between electrical load and transformer temperature. Tool Call Result: Joule’s Law is returned, indicating that heat generation is proportional to the square of the current (load), where low load leads to cooling and high load leads to heating. Response: Based on Joule’s Law, the temperature forecast should reflect the load profile. An intermediate prediction (X_int) is generated by transferring the shape of the power load curve onto the baseline temperature prediction. Tool Call: Generate intermediate prediction using shape transfer from ‘power_load‘. Tool Call Result: X_int is generated successfully, with its curve shape matching the power load pattern. Reflector: Response: Upon reviewing X_int, the temperature responds almost instantaneously to load changes. However, transformers possess significant thermal mass, and such instantaneous temperature variation violates thermal inertia. Evaluation Result: FAIL. Reason: The predicted temperature change is too abrupt and violates the heat equation. A temporal lag is required due to specific heat capacity. Refinement: A refinement is applied by introducing exponential smoothing on the rising segment ($t>60$) to simulate gradual heat accumulation caused by thermal lag. Final Output: The final forecast (X_final) is generated after refinement and is verified to satisfy physical constraints.
B.5 Reflection Mechanisms
Figure 8: Prediction performance of different reflection mechanisms across datasets.
To evaluate the impact of the reflection strategy on reasoning and forecasting quality, we compared our standard full model with an enhanced reflection variant. The latter retains the same underlying backbone LLM but introduces a more complex and iterative reflection process, along with secondary corrections based on historical segments retrieved via timestamp alignment. The results (see Figure 8) clearly show that the single-pass, continuous reasoning model achieves better performance than the enhanced reflection approach. Across both the PJM and Sunny datasets, enhanced reflection leads to significantly higher forecasting errors. This degradation stems from the disruption of temporal continuity and contextual coherence, which are crucial for modeling long-range dependencies in time series. In addition, timestamp-aligned retrieval may draw on segments generated under different external conditions, causing incorrect anchoring to seasonal patterns that do not match the current context. Moreover, iterative reflection tends to amplify minor errors, leading to compounded deviations over the full forecasting horizon. These findings suggest that maintaining a direct and uninterrupted reasoning process is essential for preserving predictive integrity.
B.6 Algorithm:CastMind Inference Pipeline
Algorithm 2 summarizes the training-free inference procedure of CastMind, including context construction, reasoning-based generation, and iterative reflective refinement with a fallback strategy.
Algorithm 2 CastMind inference pipeline (training-free).
Input: $I_{\text{in}}=(I_{\text{tp}},I_{\text{dp}},T,X_{\text{ex}},X_{\text{en}})$; toolkit $F,K,E,\mathcal{C}$.
Output: Final forecast $\hat{Y}$ (and optional explanation).
Stage 1: Context extraction.
Compute seleted features $F(X_{\text{en}})$ and select $F_{\text{sel}}\leftarrow S(F,I_{\text{in}})$.
Retrieve $(Y_{\text{case}},X_{\text{nb}},Y_{\text{nb}})\leftarrow\textsc{CaseRetrieve}(\mathcal{C},X_{\text{en}})$.
Retrieve $(K^{\prime},E^{\prime})\leftarrow\textsc{Retrieve}(K,E,I_{\text{in}})$.
Assemble $I_{\text{con}}\leftarrow(I_{\text{in}},F_{\text{sel}},K^{\prime},E^{\prime},Y_{\text{case}},X_{\text{nb}},Y_{\text{nb}})$.
Stage 2: Reasoning-based generation.
Generate intermediate prediction $Y_{\text{int}}\leftarrow\textsc{Generate}(I_{\text{con}})$.
Stage 3: Reflective evaluation and refinement.
for $r=1$ to $R_{\max}$ do
$(\texttt{ok},\texttt{fb})\leftarrow\textsc{Evaluate}(I_{\text{con}},Y_{\text{int}})$.
if $\texttt{ok}=\texttt{true}$ then return $\hat{Y}\leftarrow Y_{\text{int}}$.
$I_{\text{con}}\leftarrow\textsc{Update}(I_{\text{con}},\texttt{fb})$.
$Y_{\text{int}}\leftarrow\textsc{Generate}(I_{\text{con}})$.
end for
B.7 LLM-based Feature Selection
In CastMind, feature selection is performed by the LLM in a training-free manner. Given the contextual input $I_{\text{in}}$, the predefined feature set $F$, and the computed feature values on the current look-back window, the LLM selects relevant features for both the endogenous and exogenous sequences:
$F_{\text{sel}}^{\text{en}}=S_{\text{LLM}}\!\left(F,\phi(X_{\text{en}}),I_{\text{in}}\right),\quad F_{\text{sel}}^{\text{ex}}=S_{\text{LLM}}\!\left(F,\phi(X_{\text{ex}}),I_{\text{in}}\right),$ (5)
where $F_{\text{sel}}^{\text{en}}$ and $F_{\text{sel}}^{\text{ex}}$ denote the selected feature subsets for $X_{\text{en}}$ and $X_{\text{ex}}$, respectively. We then form the overall selected feature set as
$F_{\text{sel}}=F_{\text{sel}}^{\text{en}}\cup F_{\text{sel}}^{\text{ex}}.$ (6)
The selection is based on the forecasting requirements (e.g., horizon), variable semantics, timestamps, and the observed characteristics of the input sequences. To ensure consistent downstream usage, the LLM is instructed to output exactly $k_{\text{en}}$ feature names for $X_{\text{en}}$ and $k_{\text{ex}}$ feature names for $X_{\text{ex}}$ from $F$.
B.8 Prompt Templates for the Three-Stage Workflow
Paper-facing vs. code-facing prompts.
For clarity and reproducibility, we provide prompt templates in the paper that describe the behavior, inputs, and outputs of each stage in a general form that does not depend on specific tools. The released code also includes instructions specific to the implementation (such as function calls, logging settings, and file paths) to interact with the toolkit. These technical details do not affect the core algorithm of CastMind, and the code-facing prompts are direct implementations of the paper-facing templates based on a fixed input/output structure.
Notation-to-implementation mapping.
Table 8 maps the paper notation used throughout the manuscript to representative field names used in the code-facing prompts and intermediate artefacts.
Table 8: Mapping between paper notation and representative code-facing fields.
Paper notation Code-facing field (example)
$I_{\text{in}}$ context.input / packet.input
$I_{\text{tp}}$ task_prompt
$I_{\text{dp}}$ data_profile
$T$ timestamps
$X_{\text{en}}$ endogenous_lookback
$X_{\text{ex}}$ exogenous_lookback
$F_{\text{sel}}$ selected_features
$K$ knowledge_snippets
$E$ attributes / context_attrs
$Y_{\text{case}}$ reference_prediction
$(X_{\text{nb}},Y_{\text{nb}})$ neighbor_lookback, neighbor_pred
$I_{\text{con}}$ consolidated_context
$Y_{\text{int}}$ intermediate_prediction
Non-essential operational constraints.
The code-facing prompts include some extra rules for reliable execution (e.g., calling a tool only once, saving logs, and specifying file paths). These rules are added to make the implementation stable and the evaluation consistent. They do not change what information the model receives as input or what results are reported in the paper.
B.8.1 Stage-1 Short Template (Context Extraction)
System Prompt for Investigator Role Definition: You are the Investigator, responsible for Stage-1 context extraction in time-series forecasting. Your goal is to gather, filter, and consolidate all available information into a structured context for downstream reasoning. Input: Contextual Input: $\mathcal{I}_{\text{in}}=(\mathcal{I}_{\text{tp}},\mathcal{I}_{\text{dp}},\mathcal{T},\mathbf{X}_{\text{ex}},\mathbf{X}_{\text{en}})$ Resources: Auxiliary Toolkit (Knowledge Base, Search Engine, Memory Pool). Task Workflow: 1. Goal Alignment: Summarize the forecasting horizon $L$ and clarify variable semantics based on task profile $\mathcal{I}_{\text{tp}}$ and data profile $\mathcal{I}_{\text{dp}}$. 2. Feature Engineering: Compute feature values and select relevant temporal features for both endogenous $\mathbf{X}_{\text{en}}$ and exogenous $\mathbf{X}_{\text{ex}}$ variables to produce $\mathcal{F}_{\text{sel}}$. 3. Evidence Retrieval: Invoke the toolkit to prepare supporting evidence: Domain Knowledge $\mathcal{K}$ (e.g., physical laws). Contextual Attributes $\mathcal{E}$ (e.g., calendar events). Analogous References $(\mathbf{Y}_{\text{case}},\mathbf{X}_{\text{nb}},\mathbf{Y}_{\text{nb}})$ via retrieval. Output Format: Output a single JSON object constructing the consolidated context: $\mathcal{I}_{\text{con}}=(\mathcal{I}_{\text{in}},\mathcal{F}_{\text{sel}},\mathcal{K},\mathcal{E},\mathbf{Y}_{\text{case}},\mathbf{X}_{\text{nb}},\mathbf{Y}_{\text{nb}})$. Constraints: - Truthfulness: Do not fabricate information. If a field is unavailable, explicitly return a warning explaining the missing data. - Temporal Causality: Strictly prevent data leakage. Access to future information is restricted solely to known exogenous covariates (e.g., weather forecasts, scheduled events). Future values of the endogenous target variable are strictly prohibited.
B.8.2 Stage-2 Short Template (Reasoning-based Generation)
System Prompt for Generator Role Definition: You are the Generator, responsible for Stage-2 reasoning-based generation. Your goal is to transform the consolidated context into a refined intermediate forecast. Input: Consolidated Context: $\mathcal{I}_{\text{con}}=(\mathcal{I}_{\text{in}},\mathcal{F}_{\text{sel}},\mathcal{K},\mathcal{E},\mathbf{Y}_{\text{case}},\mathbf{X}_{\text{nb}},\mathbf{Y}_{\text{nb}})$. Task Workflow: 1. Initialization: Adopt the retrieved case $\mathbf{Y}_{\text{case}}$ as the baseline forecast. 2. Reasoning-driven Adjustment: Execute Chain-of-Thought (CoT) reasoning to analyze evidence from selected features $\mathcal{F}_{\text{sel}}$, domain knowledge $\mathcal{K}$, and neighbor references $\mathbf{Y}_{\text{nb}}$. 3. Modification: Apply targeted adjustments to the baseline only when clearly justified by the evidence; otherwise, maintain the retrieved profile. 4. Generation: Produce the intermediate forecast $\mathbf{Y}_{\text{int}}$ aligned with timestamps $\mathcal{T}$. Output Format: Output a structured object containing $\mathbf{Y}_{\text{int}}$ and a Rationale Summary (2–5 bullet points) describing the specific adjustments made and the supporting evidence used. Constraints: - Contextual Fidelity: Do not fabricate context beyond $\mathcal{I}_{\text{con}}$. - Output Efficiency: While you must perform CoT reasoning internally to ensure accuracy, do NOT output the full verbose thought process. Provide only a concise, verifiable rationale summary to facilitate efficient downstream validation.
B.8.3 Stage-3 Short Template (Reflective Evaluation and Refinement)
System Prompt for Reflector Role: Stage-3 Reflector. Validate the intermediate forecast ($\mathbf{Y}_{\text{int}}$) against evidence and physical constraints. Input: Context $\mathcal{I}_{\text{con}}$, $\mathbf{Y}_{\text{int}}$, and Rationale Summary. Task Workflow: 1. Sanity Check: Verify format, length $L$, numeric validity, and timestamp alignment with $\mathcal{T}$. 2. Critique: Check if adjustments to $\mathbf{Y}_{\text{case}}$ are solidly supported by evidence ($\mathcal{F}_{\text{sel}},\mathcal{K}$). Flag hallucinations or physical violations (e.g., mass conservation issues). 3. Refinement: If valid issues exist, execute Critique-and-Refine: propose specific corrections (e.g., time-shifting, smoothing) and generate $\mathbf{Y}_{\text{ref}}$. 4. Finalization: Output the final forecast if robust; otherwise signal failure. Output Format: Structured object with Status (PASS/FAIL), critique reasoning, and $\mathbf{Y}_{\text{final}}$. Constraints: Strictly enforce domain knowledge $\mathcal{K}$. Unlike the Generator, you MUST explicitly articulate the reasoning for any rejection or modification to ensure groundedness.

### v3

@@
-CastMind: An Interaction-Driven Agentic Reasoning Framework for Cognition-Inspired Time Series Forecasting
+AlphaCast: An Interaction-Driven Agentic Reasoning Framework for Cognition-Inspired Time Series Forecasting
 Abstract
@@
 In contrast, human experts form predictions through iterative reasoning that integrates temporal features, domain knowledge, case-based references, and supplementary context, with continuous refinement.
-In this work, we propose CastMind, an interaction-driven agentic reasoning framework that enables accurate time series forecasting with training-free large language models.
-CastMind reformulates forecasting as an expert-like process and organizes it into a multi-stage workflow involving context preparation, reasoning-based generation, and reflective evaluation, transforming forecasting from a single-pass output into a multi-turn, autonomous interaction process.
+In this work, we propose AlphaCast, an interaction-driven agentic reasoning framework that enables accurate time series forecasting with training-free large language models.
+AlphaCast reformulates forecasting as an expert-like process and organizes it into a multi-stage workflow involving context preparation, reasoning-based generation, and reflective evaluation, transforming forecasting from a single-pass output into a multi-turn, autonomous interaction process.
 To support diverse perspectives commonly considered by human experts, we develop a lightweight toolkit comprising a feature set, a knowledge base, a case library, and a contextual pool that provides external support for LLM-based reasoning.
-Extensive experiments across multiple benchmarks show that CastMind generally outperforms representative baselines [^1] https://github.com/SkyeGT/CastMind.
+Extensive experiments across multiple benchmarks show that AlphaCast generally outperforms representative baselines [^1] https://github.com/SkyeGT/CastMind.
 Machine Learning, ICML
 1 Introduction
-Figure 1: Illustration of the motivation behind CastMind: shifting forecasting from static, single-pass regression to human-like, interactive, multi-turn reasoning.
+Figure 1: Illustration of the motivation behind AlphaCast: shifting forecasting from static, single-pass regression to human-like, interactive, multi-turn reasoning.
 Time series forecasting plays a crucial role in decision-making across various real-world applications, such as energy (Cao et al., 2023), climate (Wang et al., ), and industry (Ma et al., 2025).
-Despite advances with statistical (Webby & O’Connor, 1996), deep learning (Benidis et al., 2022), and recent foundation models (Das et al., 2024), these methods typically treat forecasting as a static, single-pass regression task, generating predictions based solely on historical observations (Brockwell & Davis, 2002).
-This approach neglects reasoning, interaction, and iterative refinement, which are critical in real-world decision-making.
+Despite substantial advances with statistical (Webby & O’Connor, 1996), deep learning (Benidis et al., 2022), and recent foundation models (Das et al., 2024), these methods typically treat forecasting as a static, single-pass regression task, generating predictions based solely on historical observations (Brockwell & Davis, 2002).
+This approach neglects reasoning, interaction, and iterative refinement, which are critical in complex real-world decision-making scenarios (Cheng et al., 2026).
 In contrast, human experts rarely follow this single-pass model.
@@
 Rather than a single-pass prediction, forecasting becomes a multi-turn process where an agent actively prepares contextual information, reasons over diverse sources, and refines predictions through self-reflection.
-To operationalize this perspective, We introduce CastMind, an interaction-driven agentic reasoning framework for time series forecasting.
-CastMind leverages training-free LLMs as reasoning engines and organizes forecasting into a structured workflow: context preparation, reasoning-based generation, and reflective verification.
-By modeling interaction and refinement, CastMind bridges the gap between existing models and expert-like forecasting processes.
-CastMind implements this through a congnitive-inspired workflow that separates reasoning from prediction.
+To operationalize this perspective, We introduce AlphaCast, an interaction-driven agentic reasoning framework for time series forecasting.
+AlphaCast leverages training-free LLMs as reasoning engines and organizes forecasting into a structured workflow: context preparation, reasoning-based generation, and reflective verification.
+By modeling interaction and refinement, AlphaCast bridges the gap between existing models and expert-like forecasting processes.
+AlphaCast implements this through a congnitive-inspired workflow that separates reasoning from prediction.
 It starts with context preparation, gathering relevant information beyond raw time series, including temporal features, domain knowledge, case-based references and supplementary context.
@@
 A reflective verification stage follows, ensuring consistency and plausibility, enabling self-correction when necessary.
-To support this process, CastMind uses a lightweight toolkit with core components: a feature set, knowledge base, case library,and contextual pool.
-Together, the workflow and toolkit enable CastMind to conduct forecasting as an interactive, multi-turn reasoning process that closely mirrors expert practice.
+To support this process, AlphaCast uses a lightweight toolkit with core components: a feature set, knowledge base, case library,and contextual pool.
+Together, the workflow and toolkit enable AlphaCast to conduct forecasting as an interactive, multi-turn reasoning process that closely mirrors expert practice.
 The contributions of this work are as follows:
   • We reformulate time series forecasting from a static, single-pass regression task into an interaction-driven agentic reasoning problem, highlighting the role of iterative reasoning and refinement in forecasting.
-  • We propose CastMind, a training-free agentic reasoning framework that integrates a structured workflow with a lightweight forecasting toolkit for context-aware and reflective forecasting.
-  • Extensive experiments across multiple benchmarks show CastMind consistently outperforms statistical, deep learning, and foundation-model baselines, demonstrating its effectiveness.
+  • We propose AlphaCast, a training-free agentic reasoning framework that integrates a structured workflow with a lightweight forecasting toolkit for context-aware and reflective forecasting.
+  • Extensive experiments across multiple benchmarks show AlphaCast consistently outperforms statistical, deep learning, and foundation-model baselines, demonstrating its effectiveness.
 2 Related Work
@@
 Recent advancements in LLMs have demonstrated remarkable generalization capabilities in reasoning and decision-making (Guo et al., 2025).
-Foundational is in-context learning (ICL), enabling models to adapt to unseen tasks via prompts without parameter updates (Brown et al., 2020; Xue & Salim, 2023).
+Foundational is in-context learning (ICL), enabling models to adapt to unseen tasks via prompts without parameter updates (Xue & Salim, 2023).
 To elicit multi-step inference, prompting strategies like chain-of-thought (CoT) (Wei et al., 2022) decompose complex problems into intermediate reasoning steps.
 Subsequent research enhanced robustness through aggregation like self-consistency (Wang et al., 2022) and decomposition like least-to-most prompting (Zhou et al., 2022).
-Moving beyond static reasoning, a significant paradigm shift models LLMs as autonomous agents operating in multi-step, goal-directed settings.
+Moving beyond static reasoning, a significant paradigm shift models LLMs as autonomous agents operating in multi-step, goal-directed settings (Tao et al., 2026).
 Unlike passive inference, agentic systems maintain intermediate states and interleave reasoning with actions, invoking tools to ground outputs.
@@
 This modular design enables refinement and adaptive execution, forming the basis for our framework.
-Figure 2: Overall architecture of CastMind.
+Figure 2: Overall architecture of AlphaCast.
 The three-stage process (Investigator, Generator, Reflector) bridges human wisdom and LLM intelligence, simulating expert cognition to transform forecasting into an interactive, multi-turn reasoning process.
-3 The Proposed CastMind
-In this section, we introduce CastMind, which reformulates forecasting as a multi-stage process that simulates human-like reasoning to prepare, generate, and evaluate predictions.
+3 The Proposed AlphaCast
+In this section, we introduce AlphaCast, which reformulates forecasting as a multi-stage process that simulates human-like reasoning to prepare, generate, and evaluate predictions.
 3.1 Problem Definition
@@
 3.2 Overall Agentic Workflow
-As shown in Figure 2, CastMind is a multi-stage interaction-driven agentic reasoning framework designed for cognition-inspired time series forecasting.
+As shown in Figure 2, AlphaCast is a multi-stage interaction-driven agentic reasoning framework designed for cognition-inspired time series forecasting.
 It reformulates forecasting from a static, single-pass mapping task into an interactive, multi-turn iterative process with training-free LLM reasoning.
-Before the workflow begins, CastMind prepares a specialized toolkit comprising a feature set, a knowledge base, a case library, a contextual pool.
+Before the workflow begins, AlphaCast prepares a specialized toolkit comprising a feature set, a knowledge base, a case library, and a contextual pool.
 The process then unfolds in three coordinated stages.
-Initially, CastMind organizes useful contextual input and information retrieved from the toolkit to construct a well-defined forecasting context.
+Initially, AlphaCast organizes useful contextual input and information retrieved from the toolkit to construct a well-defined forecasting context.
 Subsequently, the framework operates on this foundation to generate reasoning-based predictions.
@@
 Feature Set.
-CastMind extracts temporal features from endogenous time series and exogenous variables to summarize temporal dynamics.
+AlphaCast extracts temporal features from endogenous time series and exogenous variables to summarize temporal dynamics.
 Given a group of extraction functions $\mathcal{F}=\{f_{1},\dots,f_{n}\}$, the feature set $F$ is constructed by applying each function to all input series:
@@
 Knowledge Base.
-CastMind maintains a knowledge base that provides prior information to support reasoning and decision-making.
+AlphaCast maintains a knowledge base that provides prior information to support reasoning and decision-making.
 The knowledge base contains: (1) practical rules, which are experience-based guidelines derived from empirical observations.
@@
 Case Library.
-To simulate the expert use of historical experiences, CastMind constructs a case library by associating each training sample $\langle X_{i},Y_{i}\rangle$ with its optimal candidate model $M_{i}$ selected from the candidate pool $\mathcal{M}$ (detailed in Appendix B.2).
+To simulate the expert use of historical experiences, AlphaCast constructs a case library by associating each training sample $\langle X_{i},Y_{i}\rangle$ with its optimal candidate model $M_{i}$ selected from the candidate pool $\mathcal{M}$ (detailed in Appendix B.2).
 To enable efficient retrieval and generation of case-based predictions, we perform K-means clustering (detailed in Appendix A.2) on all $X_{i}$, grouping them into multiple clusters.
@@
 Accurate forecasting requires a well-defined prediction task, a proper understanding of data semantics, and the availability of relevant contextual information.
-Therefore, CastMind constructs a standardized contextual input $I_{\text{in}}$ that integrates time series with textual descriptions.
+Therefore, AlphaCast constructs a standardized contextual input $I_{\text{in}}$ that integrates time series with textual descriptions.
 Specifically, $I_{\text{in}}$ comprises the task prompt, the data profile (detailed meanings of the variables), original timestamp information (Wang et al., 2024a; Zeng et al., 2024), and the raw time series $(X_{\text{ex}},X_{\text{en}})$ in original space.
@@
 3.5 The Generator: Reasoning-based Forecasting
-Given the consolidated context $I_{\text{con}}$, CastMind derives an intermediate prediction $Y_{\text{int}}$ through reasoning-driven generation instead of single-pass regression.
+Given the consolidated context $I_{\text{con}}$, AlphaCast derives an intermediate prediction $Y_{\text{int}}$ through reasoning-driven generation instead of single-pass regression.
 The model initializes with the case-based prediction $Y_{\text{case}}$ as a reliable reference.
-To handle distribution shifts between neighbor $X_{\text{nb}}$ and the current sequence $X_{\text{en}}$, CastMind uses the neighbor trajectory $Y_{\text{nb}}$ to adjust for scale and fluctuation differences (Liu et al., 2025c).
+To handle distribution shifts between neighbor $X_{\text{nb}}$ and the current sequence $X_{\text{en}}$, AlphaCast uses the neighbor trajectory $Y_{\text{nb}}$ to adjust for scale and fluctuation differences (Liu et al., 2025c).
 The framework also incorporates temporal features from the selected feature set $F_{\text{sel}}$, along with domain knowledge $K$ and supplementary context $S$.
@@
 CES 808.608 315.576 1084.713 40.776 38.327 24.218 4.124 9364.980 78.713 6.399
-CastMind 536.454 193.829 721.023 27.113 25.151 7.641 2.414 1548.825 13.294 4.771
+AlphaCast 536.454 193.829 721.023 27.113 25.151 7.641 2.414 1548.825 13.294 4.771
 MAE Sundial 10.997 10.707 8.536 3.362 3.987 2.314 1.258 31.218 6.687 1.422
@@
 CES 14.414 11.451 12.144 3.993 4.467 3.646 1.483 42.040 4.555 1.470
-CastMind 9.612 9.925 7.133 3.145 3.596 2.017 1.057 24.540 1.843 1.353
+AlphaCast 9.612 9.925 7.133 3.145 3.596 2.017 1.057 24.540 1.843 1.353
 3.6 The Reflector: Iterative Refinement
 In real-world forecasting, predictions are usually reviewed and refined before being finalized.
-Accordingly, CastMind includes a reflection stage to verify the prediction and support iterative refinement.
+Accordingly, AlphaCast includes a reflection stage to verify the prediction and support iterative refinement.
 This stage improves the overall reliability and reproducibility of the framework and consists of two parts: (1) prediction evaluation, which assesses whether the prediction satisfies the task requirements and is supported by the available evidence, ensuring that the prediction is reasonable and reliable; and (2) reasoning-trace evaluation, which examines the generated reasoning trace to ensure that each step is valid and consistent, and identifies potential issues such as unsupported claims or logical gaps.
 Together, these checks help ensure that both the forecast and its underlying reasoning adhere to the task requirements.
-If issues are found, CastMind applies targeted, evidence-supported corrections to $Y_{\text{int}}$ to obtain a refined prediction.
+If issues are found, AlphaCast applies targeted, evidence-supported corrections to $Y_{\text{int}}$ to obtain a refined prediction.
 The refined result is then re-evaluated with the same criteria.
-If acceptable, CastMind outputs the forecast; otherwise, it returns feedback to the previous stage for another update, repeating until an acceptable result is obtained.
+If acceptable, AlphaCast outputs the forecast; otherwise, it returns feedback to the previous stage for another update, repeating until an acceptable result is obtained.
 4 Experiments
-In this section, we conduct experiments on various forecasting benchmarks comparing CastMind with mainstream approaches to demonstrate its effectiveness.
+In this section, we conduct experiments on various forecasting benchmarks comparing AlphaCast with mainstream approaches to demonstrate its effectiveness.
 4.1 Experimental Setup
@@
 4.2 Main Results
-Table 1 compares CastMind against twelve baselines across short-term and long-term benchmarks.
-Overall, CastMind achieves state-of-the-art performance with superior accuracy across most metrics.
-In short-term scenarios, CastMind shows outstanding stability in the face of high volatility, effectively managing unpredictable fluctuations.
-Unlike statistical methods such as ARIMA, which struggle with rapid non-linear fluctuations and inherent stochastic volatility in price dynamics, or deep learning models such as PatchTST, which may over-smooth dynamics, CastMind effectively captures sharp peaks by leveraging historical references to avoid phase shifts.
+Table 1 compares AlphaCast against twelve baselines across short-term and long-term benchmarks.
+Overall, AlphaCast achieves state-of-the-art performance with superior accuracy across most metrics.
+In short-term scenarios, AlphaCast shows outstanding stability in the face of high volatility, effectively managing unpredictable fluctuations.
+Unlike statistical methods such as ARIMA, which struggle with rapid non-linear fluctuations and inherent stochastic volatility in price dynamics, or deep learning models such as PatchTST, which may over-smooth dynamics, AlphaCast effectively captures sharp peaks by leveraging historical references to avoid phase shifts.
 This gives it a significant advantage on datasets like BE and FR.
-In long-term settings, CastMind exhibits superior forecasting capability, especially on datasets like Windy Power and Sunny Power, which exhibit complex environmental dependencies.
-While foundation models such as Chronos perform competitively due to their zero-shot generalization, CastMind maintains a consistent lead through its reasoning and refinement capabilities that enforce causal constraints derived from diverse contexts.
-For instance, on the Windy Power dataset, CastMind substantially outperforms baselines like DLinear.
+In long-term settings, AlphaCast exhibits superior forecasting capability, especially on datasets like Windy Power and Sunny Power, which exhibit complex environmental dependencies.
+While foundation models such as Chronos perform competitively due to their zero-shot generalization, AlphaCast maintains a consistent lead through its reasoning and refinement capabilities that enforce causal constraints derived from diverse contexts.
+For instance, on the Windy Power dataset, AlphaCast substantially outperforms baselines like DLinear.
 This superiority across different scales validates the effectiveness of the interactive, multi-turn reasoning process in handling diverse temporal patterns and ensuring forecasting stability.
@@
 4.3 Ablation studies
-This section validates the CastMind framework through a multi-perspective ablation study, covering both architectural stages, contextual input, and auxiliary toolkit modules.
+This section validates the AlphaCast framework through a multi-perspective ablation study, covering both architectural stages, contextual input, and auxiliary toolkit modules.
 Ablation Study on Framework Components.
@@
 As shown in Table 2, the full model consistently achieves best performance, validating component collaboration.
-Pecifically, removing the toolkit yields highest errors in scenarios like BE and ETTm, confirming external context is foundational.
+Specifically, removing the toolkit yields highest errors in scenarios like BE and ETTm, confirming external context is foundational.
 Furthermore, comparing the baseline lacking reasoning, which serves as the state-of-the-art candidate model backbone, with the full model reveals that reasoning-based forecasting substantially improves accuracy, evidenced by notable error reduction on the FR dataset.
@@
 To verify the necessity of constructing a comprehensive context $I_{\text{con}}$, we conduct a component-wise ablation study.
-As shown in Figure 3, the complete CastMind framework consistently yields the lowest MAE, confirming the synergistic contribution of all elements.
+As shown in Figure 3, the complete AlphaCast framework consistently yields the lowest MAE, confirming the synergistic contribution of all elements.
 Specifically, removing timestamps causes significant degradation on periodic datasets such as PJM and Sunny Power, highlighting the critical role of explicit temporal identifiers in capturing seasonality that numerical sequences alone cannot convey.
@@
 4.4 Exploration Analysis
-Figure 6: Feature-level and model-level case analysis of CastMind across datasets.
+Figure 6: Feature-level and model-level case analysis of AlphaCast across datasets.
 The heatmaps report the relative usage frequency (%) of temporal features and candidate forecasting models, highlighting a preference for statistical features and deep learning models.
-Figure 7: Case-study comparisons on four datasets (ETTm, MOPEX, Windy Power, and DE), where CastMind produces predictions that track the ground truth more closely than the cased-based prediction, especially under periodic or highly volatile patterns.
+Figure 7: Case-study comparisons on four datasets (ETTm, MOPEX, Windy Power, and DE), where AlphaCast produces predictions that track the ground truth more closely than the cased-based prediction, especially under periodic or highly volatile patterns.
 Table 4: Impact of generation strategies. Holistic generation outperforms the step-wise baseline, demonstrating the necessity of coherent, one-shot prediction.
@@
 As illustrated in Figure 4, GPT-5 consistently achieves the lowest MAE across diverse datasets such as Windy Power, DE, and PJM, establishing a clear performance advantage over other LLMs.
-The superior performance of GPT-5 is attributed to its powerful pre-training and reasoning capabilities, enabling it to effectively handle complex time series, adapt to various context, and provide more accurate predictions.
+The superior performance of GPT-5 is attributed to its powerful pre-training and reasoning capabilities, enabling it to effectively handle complex time series, adapt to various contexts, and provide more accurate predictions.
 Notably, the performance gap becomes most distinct on the Windy Power dataset, where GPT-5 demonstrates a significantly stronger ability to model complex environmental dependencies and high-frequency volatility compared to competitive baselines.
@@
 Preference Analysis on Toolkit.
-We analyze the usage frequency within the feature set and case library to understand CastMind’s preferences.
+We analyze the usage frequency within the feature set and case library to understand AlphaCast’s preferences.
 As shown in Figure 6, heatmaps reveal selection patterns driven by data characteristics.
-In the feature domain, CastMind favors fundamental statistical features, suggesting that reliable global properties serve as better foundations than volatile complexity measures.
+In the feature domain, AlphaCast favors fundamental statistical features, suggesting that reliable global properties serve as better foundations than volatile complexity measures.
 Meanwhile, model selection shows a clear bias towards deep learning architectures, with Transformers consistently dominating the candidate pool over statistical models, confirming that high-capacity backbones provide the necessary prior.
-Ultimately, the varying distribution across datasets confirms that the framework dynamically adjusts its retrieval strategy, rather than relying on static priors.
+Ultimately, the varying distribution across datasets confirms that the framework dynamically adjusts its retrieval strategy, rather than relying on priors.
 Case-based Prediction Analysis.
@@
 As shown in Figure 7, the case-based prediction often suffers from over-smoothing, generating conservative trajectories that miss rapid fluctuations.
-This is evident on MOPEX and Windy Power, where the candidate models misses significant peaks and reverts to the mean.
-In contrast, CastMind shows superior sensitivity, closely tracking the ground truth even under high volatility on the ETTm and DE.
+This is evident on MOPEX and Windy Power, where the candidate models miss significant peaks and reverts to the mean.
+In contrast, AlphaCast shows superior sensitivity, closely tracking the ground truth even under high volatility on the ETTm and DE.
 This confirms that reasoning allows the model to escape the mean-reversion trap, ensuring forecasts retain the original series’ characteristics.
@@
 5 Conclusion
-In this work, we introduces CastMind, an interaction-driven agentic reasoning framework designed to bridge the gap between existing time series forecasting models and expert-driven iterative reasoning.
-CastMind redefines the forecasting process as a multi-turn iteration, integrating context preparation, reasoning-based generation, and reflective evaluation, transforming forecasting from a single-pass output into a multi-turn autonomous interaction process.
+In this work, we introduce AlphaCast, an interaction-driven agentic reasoning framework designed to bridge the gap between existing time series forecasting models and expert-driven iterative reasoning.
+AlphaCast redefines the forecasting process as a multi-turn iteration, integrating context preparation, reasoning-based generation, and reflective evaluation, transforming forecasting from a single-pass output into a multi-turn autonomous interaction process.
 The proposed framework leverages training-free LLMs as reasoning engines, complemented by a lightweight toolkit that effectively combines human wisdom with LLM intelligence.
-Comprehensive evaluations demonstrate that CastMind’s ability to model expert-like iterative reasoning and refinement is key to its superior performance in both short-term and long-term forecasting tasks.
+Comprehensive evaluations demonstrate that AlphaCast’s ability to model expert-like iterative reasoning and refinement is key to its superior performance in both short-term and long-term forecasting tasks.
 Although the results are preliminary, our findings suggest that forecasting is shifting from a static regression-based approach to a interaction-driven agentic reasoning paradigm.
 Impact Statement
-This work advances time series forecasting through CastMind, an interaction-driven agentic framework that enhances transparency and decision reliability by mimicking expert-like iterative reasoning.
+This work advances time series forecasting through AlphaCast, an interaction-driven agentic framework that enhances transparency and decision reliability by mimicking expert-like iterative reasoning.
 By transforming forecasting into a multi-stage, autonomous interaction process supported by external context, our approach significantly improves interpretability in critical sectors such as infrastructure management and financial planning.
@@
 Lastly, seasonal strength measures the dominance of seasonal patterns relative to residual variation (Wang et al., 2006), guiding the model to adaptively handle strong or weak seasonality.Collectively, this comprehensive temporal feature set equips CastMind with a deeper understanding of the time series, supporting effective reasoning and self-correction.
-Table 5: Mathematical definitions of the twenty extracted temporal features used in CastMind, covering distributional statistics, temporal dependence, spectral complexity, structural heterogeneity, and seasonality, serving as a compact summary of key temporal characteristics for forecasting.
+Table 5: Mathematical definitions of the twenty extracted temporal features used in AlphaCast, covering distributional statistics, temporal dependence, spectral complexity, structural heterogeneity, and seasonality, serving as a compact summary of key temporal characteristics for forecasting.
 Feature Description Formula / Mathematical Definition
@@
 Table 7: Forecasting performance on short-term and long-term benchmarks.
-We compare CastMind with each individual candidate model in the case library to show the benefit of interaction-driven reasoning and iterative refinement.
+We compare AlphaCast with each individual candidate model in the case library to show the benefit of interaction-driven reasoning and iterative refinement.
 Setting Short-term Forecasting Long-term Forecasting
 Models Metrics BE DE FR NP PJM ETTh ETTm Windy Power Sunny Power MOPEX
-CastMind MSE 536.454 193.829 721.023 27.113 25.151 7.641 2.414 1548.825 13.294 4.771
+AlphaCast MSE 536.454 193.829 721.023 27.113 25.151 7.641 2.414 1548.825 13.294 4.771
 MAE 9.612 9.925 7.133 3.145 3.596 2.017 1.057 24.540 1.843 1.353